# PHENIX Collaboration

### v1

@@ VII Comparison with Monte-Carlo simulations
The low-$m_{T}$ decrease of the $\lambda(m_{T})$ correlation-strength measurements [19, 39, 20] of charged pions is found in Ref. [34] to be an indirect signal of the in-medium mass reduction of the $\eta^{\prime}$ particles. The $\lambda(m_{T})$ suppression at low momenta has been investigated in Au$+$Au collisions at $\sqrt{s_{NN}}$= 200 GeV collisions by Monte-Carlo simulations [40, 41]. These calculations were also compared to the results of the previous PHENIX analysis of 0%–30% central Au$+$Au collisions at $\sqrt{s_{NN}}=200$ GeV [1]. Although the measurements are performed in pair-$m_{T}$ (cf. Section VI) ranges while the simulations are in the single-particle transverse mass $\sqrt{m^{2}+p_{T}^{2}}$, this difference can be neglected as $\lambda$ is a characteristic parameter of the two-particle correlation function at vanishing values of relative momentum, where $K{=}p$ and hence the single-particle transverse mass is equal to the pair $m_{T}$. This property is also utilized in the core-halo picture, where the smoothness approximation is warranted and $\lambda$ is evaluated at $K{=}p$ [34].
The comparison of Monte-Carlo resonance-model simulations of $\lambda(m_{T})/\lambda_{\rm max}$ to the above presented data shows that, within systematic uncertainties, an in-medium mass drop of $\eta^{\prime}$ is not inconsistent with our measured data. Theoretically, the mass of the $\eta^{\prime}$ meson could be sensitive to the $U_{A}(1)$ symmetry restoration in hot and dense hadronic matter [42, 43]. The key point being that compared to the other pseudoscalar mesons, the $\eta^{\prime}$ is anomalously heavy, $\approx$958 MeV, although the $\eta^{\prime}$ quark content is similar to that of the $\eta$ meson. The $\eta$ meson mass is closer to that of the kaon, $\approx$548 MeV, which is about 410 MeV lighter than that of the $\eta^{\prime}$ [44]. The extra mass is explained in the Standard Model in terms of the $U_{A}(1)$ anomaly that couples the mass of the $\eta^{\prime}$ to the topological properties of the quantum-chromodynamics vacuum state. If at high temperatures the structure of this vacuum changes, the extra 410 MeV mass difference may vanish and the $\eta^{\prime}$ mesons may return to the mass scale of the other pseudoscalar mesons, with its mass becoming similar to that of the $\eta$ [43]. As the $\eta^{\prime}$ meson leaves the hot and dense matter, it regains its anomalously large mass at the expense of its kinetic energy and consequently this effect modifies the spectrum. See below for a discussion of how Monte-Carlo calculations are used to investigate the effect on $\lambda(m_{T})$ of this mass modification, along the lines of Ref. [1].
@@
Figure 11: Comparison of the data to the optimal fits of the six centrality classes and the case where the mass of the $\eta^{\prime}$ meson in hot hadronic matter is equal to the mass of the $\eta^{\prime}$ meson in vacuum.
Our results suggest that a significant, centrality independent, in-medium mass drop of the $\eta^{\prime}$ is not inconsistent with the present measurements. In Figs. 12 and 13 the optimal values of $B^{-1}_{\eta^{\prime}}$ and $m^{*}_{\eta^{\prime}}$, respectively, are shown as function of $N_{\rm part}$. The average value of the modified $\eta^{\prime}$ mass is $m^{*}_{\eta^{\prime}}=581^{+12}_{-20}{\rm(stat)}^{+205}_{-91}{\rm(syst)}$ MeV. In the case of the inverse slope parameter, the point corresponding to the 0%–10% centrality bin was omitted from the constant fit, i.e, the average value indicated in Fig. 12, represents the average value of the five remaining points, that is $B_{\eta^{\prime}}^{-1}=56^{+22}_{-14}{\rm(stat)}^{+190}_{-31}{\rm(syst)}$ MeV. In Figs. 12 and 13 the dashed lines denote these centrality-averaged values. In at least five centrality classes, from 0%–10% to 40%–50%, the measured $\lambda(m_{T})/\lambda_{\rm max}$ functions are found to be consistently described by considering the suppressed mass of the $\eta^{\prime}$. However, in the most peripheral centrality class (50%–60%) an unmodified $\eta^{\prime}$ mass cannot be excluded. This observation can be interpreted as the lack of mass modification due to the lack of enough hot and dense matter. The data are not inconsistent with Monte-Carlo simulations with vanishing differences between $m^{*}_{\eta^{\prime}}$ and $m_{\eta}{\approx}548$ MeV.
@@
Figure 13: The centrality dependence of the best values of the in-medium mass of the $m^{*}_{\eta^{\prime}}$ are shown with symbols, together with statistical (bars) and systematic (red boxes) uncertainties. The average (using only statistical uncertainties) of the values of $m^{*}_{\eta^{\prime}}$ found in the six centrality intervals is indicated by a dashed (red) line. The fitted values are compared to theoretical predictions of Weinberg [55], Horvatić, Kekez and Klabučar  [56], Pisarski and Wilczek [42], Kapusta, Kharzeev and McLerran [43], Huang and Wang [57], and Kwon, Lee, Morita, and Wolf [58].
@@
[…]
  • Recent calculations by Horvatić, Kekez and Klabučar (HKK) [56], based on the calculations of Witten-Veneziano equation and the generalization by Shore, who evaluated the properties of the $\eta$ and $\eta^{\prime}$ mesons at high temperatures, when the $U_{A}(1)$ and the chiral symmetry breaking are considered together. A substantial decrease in the $\eta^{\prime}$ mass around the chiral transition temperature was obtained in Ref. [56], but there was no decrease in the $\eta$ mass, which improves on earlier results in Ref. [40, 41] that are consistent within uncertainties. The lower limits, which are shown by a light-green dashed-dotted line in Fig. 13, lie above our results except in the 50%–60% centrality class, where the uncertainty on our value of $m^{*}_{\eta^{\prime}}$ is particularly large. This is also true of the similar limit of Huang and Wang [57]. These models are in a modest tension with our results.
  • The Pisarski-Wilczek lower limit of 600 MeV, as determined from Fig. 1 of Ref. [42], is shown as a solid black line. It is, within uncertainty, consistent, with the fitted values.
  • The significant in-medium mass drop of the $\eta^{\prime}$ meson was related to the restoration of $U_{A}(1)$ symmetry and described as the return of a Goldstone-boson by Kapusta, Kharzeev and McLerran (KKM) Ref. [43]. They have given a broad range for the possible in-medium mass of the $\eta^{\prime}$ meson in case of a partial ${\rm U}_{A}(1)$ symmetry restoration, namely $411$ MeV $\leq m^{*}_{\eta^{\prime}}\leq 685$ MeV shown as a box in Fig. 13. Our results lie within this range.
  • Kwon, Lee, Morita, and Wolf (KLMW) also utilized a temperature-dependent, generalized Witten–Veneziano relation to obtain a nearly 50% decrease of the mass of the $\eta^{\prime}$ meson in hot and dense hadronic medium, as a consequence of the restoration of $U_{A}(1)$ symmetry [58]. Their 2012 prediction of $m^{*}_{\eta^{\prime}}\geq 464$ MeV, shown as a triple dotted-dashed line in Fig. 13, is consistent with our results in each investigated centrality class.
@@ VIII Summary and Conclusion
This paper presents measurements of the two-pion BEC function in Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$= 200 GeV using data recorded in 2010 by the PHENIX experiment. Lévy stable distributions are utilized to characterize the data and determine the transverse-mass and centrality dependence of the Lévy parameters.
The Lévy parameterization is found to give a statistically acceptable description of the data with the Lévy exponent, $\alpha$, significantly greater than 1 and less than 2. This exponent is determined in 23 $m_{T}$ and in 6 centrality bins and is observed that to be well described with its $m_{T}$-averaged value in every centrality bin. However, these averages do depend on centrality.
The Lévy-scale parameter $R$ is proportional to the HWHM of the source distribution, with a coefficient of proportionality that depends on the Lévy exponent $\alpha$. The behavior of its inverse square as function of $m_{T}$ that is observed can be described by an affine linear fit, whose parameters are examined as a function of centrality. The parameter $A$, which can be related to the transverse velocity of the expansion, shows a trend similar to hydrodynamical predictions [36, 37]. The assumption of local thermalization in these calculations also results in the $\alpha=2$, Gaussian special case [36, 25, 37].
Within statistical uncertainties, the parameter $B$, is consistent with zero or with a slightly negative value. A negative value of $B$ is possible if the Cooper-Frye freeze-out terms are also taken into account [59]. This result may indicate a source that includes local thermalization with hydrodynamical expansion, followed by rescattering and decays of resonances. However, such nonequilibrium, scale dependent features typically would result in deviations from the Lévy shape and from the applicability of generalized central-limit theorems. But first-order deviations from the Lévy-stable source distributions using the expansion method of Refs. [23, 9, 10] were found to be consistent with zero. It is theoretically challenging to explain simultaneously the measured value of $\alpha$, which is found to be significantly less than the Gaussian value of 2, and the $m_{T}$ dependence of the Lévy scale parameter $R$, which follows a hydrodynamically predicted scaling, (see, e.g., Ref. [38]).
@@
To quantify the phenomena, the Gaussian width and amplitude parameters, $\sigma$ and $H$, were introduced. These parameters are observed to be centrality independent, except in the 50%–60% case, where the uncertainty on $H$ increases and does not allow a statistically significant conclusion. This independence is helpful to rule out or validate models with predictions about this behavior. The pion laser model, for example, predicts a strong centrality and multiplicity dependence [60, 61, 40].
In Ref. [34] it was observed that the amount of suppression of $\lambda(m_{T})$ and hence of $\lambda(m_{T})/\lambda_{\rm max}$ at low $m_{T}$ can be an experimentally observable signal of the (partial) restoration of $U_{A}(1)$ symmetry and a measure of in-medium reduction of the mass of the anomalously heavy $\eta^{\prime}$ mass. An approximately constant trend of the intercept parameter $\lambda$ as a function of $m_{T}$ was observed in small systems at lower energies, ($E_{\rm LAB}=150$ AGeV [7, 18]), which is consistent with vanishing suppression ($H=0$). These results are both qualitatively and quantitatively different from the presence of the suppression ($H=0.42\pm 0.02$), which is within uncertainties independent of the centrality in $\sqrt{s_{NN}}=200$ GeV Au$+$Au collisions at RHIC. This entirely experimental data-based, Monte-Carlo-independent observation suggests that the signal characterized by the value of $H$ is dependent on the energy and/or the system size. Further detailed measurements are needed to determine the energy and system size, where $H$ becomes nonvanishing for the first time, starting from zero in the small Be+Be collisions at the relatively low energy of $E_{\rm LAB}=150$ AGeV.
@@ Acknowledgements.
We thank the staff of the Collider-Accelerator and Physics Departments at Brookhaven National Laboratory and the staff of the other PHENIX participating institutions for their vital contributions. We acknowledge clarifying and inspiring discussions with Sándor Hegyi, Dubravko Klabučar, Robert Pisarski and György Wolf concerning the interpretation of their published results. S.L. thanks IFJ PAN, Krakow, Poland for their encouragement to finish this work. We acknowledge support from the Office of Nuclear Physics in the Office of Science of the Department of Energy, the National Science Foundation, Abilene Christian University Research Council, Research Foundation of SUNY, and Dean of the College of Arts and Sciences, Vanderbilt University (U.S.A), Ministry of Education, Culture, Sports, Science, and Technology and the Japan Society for the Promotion of Science (Japan), Conselho Nacional de Desenvolvimento Científico e Tecnológico and Fundação de Amparo à Pesquisa do Estado de São Paulo (Brazil), Natural Science Foundation of China (People’s Republic of China), Croatian Science Foundation and Ministry of Science and Education (Croatia), Ministry of Education, Youth and Sports (Czech Republic), Centre National de la Recherche Scientifique, Commissariat à l’Énergie Atomique, and Institut National de Physique Nucléaire et de Physique des Particules (France), J. Bolyai Research Scholarship, EFOP, HUN-REN ATOMKI, NKFIH, and OTKA (Hungary), Department of Atomic Energy and Department of Science and Technology (India), Israel Science Foundation (Israel), Basic Science Research and SRC(CENuM) Programs through NRF funded by the Ministry of Education and the Ministry of Science and ICT (Korea). Ministry of Education and Science, Russian Academy of Sciences, Federal Agency of Atomic Energy (Russia), VR and Wallenberg Foundation (Sweden), University of Zambia, the Government of the Republic of Zambia (Zambia), the U.S. Civilian Research and Development Foundation for the Independent States of the Former Soviet Union, the Hungarian American Enterprise Scholarship Fund, the US-Hungarian Fulbright Foundation, and the US-Israel Binational Science Foundation.
@@
Monte-Carlo simulations with scenarios that allow for (or exclude) a possible in-medium mass modification of the $\eta^{\prime}$ meson and their comparisons to data were performed in Refs. [34, 40, 41] and in the previous PHENIX paper (Ref. [1]) by utilizing standard $\chi^{2}$ and CL maps. Similar, but now centrality-dependent simulations are summarized in this Appendix. These comparisons result in $\chi^{2}$ and CL maps that are utilized to determine the $\chi^{2}$ minimum (or CL maximum) yielding the best values, as well as the statistical and the systematic uncertainties of a possible in-medium mass-drop of the $\eta^{\prime}$ meson. Figure 14 shows an example of such a comparison, where each bin corresponds to a comparison of Monte-Carlo simulations of resonance decay chains with our data on $\lambda(m_{T})/\lambda_{m{\rm max}}$ as a function of centrality, as detailed in Section VII.
Each panel of Fig. 14 contains the optimal values of $m^{*}_{\eta^{\prime}}$ and $B_{\eta^{\prime}}^{-1}$ in the scanned region. For our simulations, a unique $\chi^{2}$ minimum is found, which is the dominant mechanism for soft particle production and is known to be thermal in $\sqrt{s_{{}_{NN}}}=200$ GeV Au$+$Au collisions. Earlier, a significantly broader class of models was considered in Refs. [40, 41] that resulted in two, characteristically different minima in the 0%–30% centrality class of the same reaction. Refs. [40, 41] evaluated the exclusion limits, focusing in particular on the smallest possible in-medium mass drop of the $\eta^{\prime}$ meson, which is required to describe the data with at least 0.1% CL in any of the considered Monte-Carlo simulations.
Our systematic investigations include an estimation of the exclusion limits for the lowest- and largest-possible values of the in-medium mass of the $\eta^{\prime}$ meson. The systematic variations are consistent, i.e. all the systematic variations result in essentially the same minimum. In contrast to Refs. [40, 41], our indirect observation of in-medium $\eta^{\prime}$ mass modification is used with the best value and its statistical and systematic uncertainties of $m^{*}_{\eta^{\prime}}$, rather than exclusion limits. Table 2 shows estimates of the exclusion limits, which do not result in statistically significant new details; hence they are not reported in detail here.
@@
The $\eta^{\prime}$ spectrum has six parameters, which are considered as inputs to the $\lambda(m_{T})$ simulations. The two most important inputs are the in-medium mass of the $\eta^{\prime}$, $m^{*}_{\eta^{\prime}}$, and the characteristic-slope parameter, $B_{\eta^{\prime}}^{-1}$, of the $\eta^{\prime}$ mesons that are emitted with PDG mass after the decay of the condensate. The second parameter characterizes a second component, a low-$p_{T}$ part of the $\eta^{\prime}$ spectrum and determines the slope of the suppression of the $\lambda$ parameter at low-$p_{T}$ [34, 40, 41]. These two parameters determine the shape of the observed dip in the low-$m_{T}$ region of the $\lambda(m_{T})$ function: a strongly reduced $\eta^{\prime}$ mass causes a dip in the low-$m_{T}$ part of the $\lambda(m_{T})$ function, while $B_{\eta^{\prime}}^{-1}$ controls the width of this dip. These two parameters are considered as fit variables. The values of four nuisance parameters $\alpha_{m{\rm th}}$, $T_{f}$, $\langle u_{mT}\rangle$, and $T_{m{\rm cond}}$, are treated as constants in the fit, but are varied to evaluate the systematic uncertainties of the fitted values.
The $\chi^{2}$ scans provide fine grids of CL in each centrality class. The in-medium mass is scanned with 10 MeV steps between 0 and 958 MeV (the vacuum mass of $\eta^{\prime}$) and $B_{\eta^{\prime}}^{-1}$ with 20 MeV steps between 0 and 350 MeV. The minimum value of $\chi^{2}$ (or maximum value of CL) is determined, which provides the most likely values of $m^{*}_{\eta^{\prime}}$ and $B_{\eta^{\prime}}^{-1}$ and their statistical uncertainties, as well as the CL of the fits to the data, which are compared to the data in Figure 11. The values of the four “constant” parameters, $\alpha_{m{\rm th}}$, $T_{f}$, $\langle u_{m_{T}}\rangle$, and $T_{m{\rm cond}}$, are varied to evaluate the systematic uncertainties of the fitted values:
$\alpha_{m{\rm th}}$

### v2

@@ VII Comparison with Monte-Carlo simulations
The low-$m_{T}$ decrease of the $\lambda(m_{T})$ correlation-strength measurements [21, 42, 22] of charged pions is found in Ref. [36] to be an indirect signal of the in-medium mass reduction of the $\eta^{\prime}$ particles. The $\lambda(m_{T})$ suppression at low transverse mass ($m_{T}$) has been investigated in Au$+$Au collisions at $\sqrt{s_{NN}}$= 200 GeV collisions by Monte-Carlo simulations [43, 44]. These calculations were also compared to the results of the previous PHENIX analysis of 0%–30% central Au$+$Au collisions at $\sqrt{s_{NN}}=200$ GeV [1]. Although the measurements are performed in pair-$m_{T}$ (cf. Section VI) ranges while the simulations are in the single-particle transverse mass $\sqrt{m^{2}+p_{T}^{2}}$, this difference can be neglected as $\lambda$ is a characteristic parameter of the two-particle correlation function at vanishing values of relative momentum, where $K{=}p$ and hence the single-particle transverse mass is equal to the pair $m_{T}$. This property is also utilized in the core-halo picture, where the smoothness approximation is warranted and $\lambda$ is evaluated at $K{=}p$ [36].
The comparison of Monte-Carlo resonance-model simulations of $\lambda(m_{T})/\lambda_{\rm max}$ to the above presented data shows that, within systematic uncertainties, an in-medium mass drop of $\eta^{\prime}$ is not inconsistent with our measured data. Theoretically, the mass of the $\eta^{\prime}$ meson could be sensitive to the $U_{A}(1)$ symmetry restoration in hot and dense hadronic matter [45, 46]. The key point being that compared to the other pseudoscalar mesons, the $\eta^{\prime}$ is anomalously heavy, $\approx$958 MeV, although the $\eta^{\prime}$ quark content is similar to that of the $\eta$ meson. The $\eta$ meson mass, $m_{\eta}\approx$548 MeV, is closer to the mass of the charged kaons, $\approx{494}$ MeV while it is 410 MeV lighter than the mass of the $\eta^{\prime}$ meson, $m_{\eta}^{\prime}\approx{958}$ MeV, although the $\eta$ and the $\eta^{\prime}$ mesons have the same quark content  [47]. The extra mass is explained in the Standard Model in terms of the $U_{A}(1)$ anomaly that couples the mass of the $\eta^{\prime}$ to the topological properties of the quantum-chromodynamics vacuum state. If at high temperatures the structure of this vacuum changes, the extra 410 MeV mass difference may vanish and the $\eta^{\prime}$ mesons may return to the mass scale of the other pseudoscalar mesons, with its mass becoming similar to that of the $\eta$ [46]. As the $\eta^{\prime}$ meson leaves the hot and dense matter, it regains its anomalously large mass at the expense of its kinetic energy and consequently this effect modifies the spectrum. See below for a discussion of how Monte-Carlo calculations are used to investigate the effect on $\lambda(m_{T})$ of this mass modification, along the lines of Ref. [1, 1, 36, 43, 44].
@@
Figure 11: Comparison of the data to the optimal fits of the six centrality classes (solid line) and the case where the mass of the $\eta^{\prime}$ meson in hot hadronic matter is equal to the mass of the $\eta^{\prime}$ meson in vacuum (dashed line).
Our results suggest that a significant, centrality independent, in-medium mass drop of the $\eta^{\prime}$ is not inconsistent with the present measurements. In Figs. 12 and 13 the optimal values of $B^{-1}_{\eta^{\prime}}$ and $m^{*}_{\eta^{\prime}}$, respectively, are shown as function of $N_{\rm part}$. The average value of the modified $\eta^{\prime}$ mass is $m^{*}_{\eta^{\prime}}=581^{+12}_{-20}{\rm(stat)}^{+205}_{-91}{\rm(syst)}$ MeV. In the case of the inverse-slope parameter, the point corresponding to the 0%–10% centrality bin was omitted from the constant fit, i.e, the average value indicated in Fig. 12, represents the average value of the five remaining points, that is $B_{\eta^{\prime}}^{-1}=56^{+22}_{-14}{\rm(stat)}^{+190}_{-31}{\rm(syst)}$ MeV. In Fig. 12 the solid [black] line denotes this centrality-averaged value. In at least five centrality classes, from 0%–10% to 40%–50%, the measured $\lambda(m_{T})/\lambda_{\rm max}$ functions are found to be consistently described by considering the suppressed mass of the $\eta^{\prime}$. However, in the most-peripheral centrality class (50%–60%) an unmodified $\eta^{\prime}$ mass cannot be excluded. In Fig. 13, the dashed [red] line indicates the centrality-averaged value of the in-medium modified mass of the $\eta^{\prime}$ meson. This observation of minimal mass modification can be interpreted as being due to the lack of enough hot and dense matter. The data are not inconsistent with Monte-Carlo simulations with vanishing differences between $m^{*}_{\eta^{\prime}}$ and $m_{\eta}{\approx}548$ MeV.
@@
Figure 13: The centrality dependence of the best values of the in-medium mass of the $m^{*}_{\eta^{\prime}}$ are shown with full circle symbols, together with statistical (vertical lines) and systematic (red boxes) uncertainties. The average (using only statistical uncertainties) of the values of $m^{*}_{\eta^{\prime}}$ found in the six centrality intervals is indicated by a dashed (red) line. The fitted values are compared to theoretical predictions of Weinberg [58], Horvatić, Kekez and Klabučar  [59], Pisarski and Wilczek [45], Kapusta, Kharzeev and McLerran [46], Huang and Wang [60], and Kwon, Lee, Morita, and Wolf [61].
@@
  • The Weinberg upper limit, $m^{*}_{\eta^{\prime}}\leq\sqrt{3}m_{\pi}$, suggested in Ref. [58] clearly overestimates significantly, in each centrality bin, the possible in-medium mass drop of the $\eta^{\prime}$ particles.
  • Recent calculations by Horvatić, Kekez, and Klabučar (HKK) [59], based on the calculations of Witten-Veneziano equation and the generalization by Shore, evaluated the properties of the $\eta$ and $\eta^{\prime}$ mesons at high temperatures, when the $U_{A}(1)$ and the chiral symmetry breaking are considered together. A substantial decrease in the $\eta^{\prime}$ mass around the chiral transition temperature was obtained in Ref. [59], but there was no decrease in the $\eta$ mass. The new HKK results are an improvement on the earlier results of Ref. [43, 44]. The lower limits, which are shown by the [light-green] dashed-dotted line in Fig. 13, lie above our results except in the 50%–60% centrality class, where the uncertainty on our value of $m^{*}_{\eta^{\prime}}$ is particularly large. This is also true of the similar limit of Huang and Wang [60]. These models are in a modest tension with our results.
  • The Pisarski-Wilczek lower limit of 600 MeV, as determined from Fig. 1 of Ref. [45], is shown as a solid [black] line. It is, within uncertainty, consistent, with the fitted values.
  • The significant in-medium mass drop of the $\eta^{\prime}$ meson was related to the restoration of $U_{A}(1)$ symmetry and described as the return of a Goldstone-boson by Kapusta, Kharzeev and McLerran (KKM) Ref. [46]. They have given a broad range for the possible in-medium mass of the $\eta^{\prime}$ meson in case of a partial ${\rm U}_{A}(1)$ symmetry restoration, namely $411$ MeV $\leq m^{*}_{\eta^{\prime}}\leq 685$ MeV shown as a [green] box in Fig. 13. Our results lie within this range.
  • Kwon, Lee, Morita, and Wolf (KLMW) also utilized a temperature-dependent, generalized Witten-Veneziano relation to obtain a nearly 50% decrease in the mass of the $\eta^{\prime}$ meson in hot and dense hadronic medium, as a consequence of the restoration of $U_{A}(1)$ symmetry [61]. Their 2012 prediction of $m^{*}_{\eta^{\prime}}\geq 464$ MeV, shown as a triple dotted-dashed line in Fig. 13, is consistent with our results in each investigated centrality class.
@@ VIII Summary and Conclusion
This paper presents measurements of the two-pion BEC function in Au$+$Au collisions at $\sqrt{s_{{}_{NN}}}$= 200 GeV using data recorded in 2010 by the PHENIX experiment. Lévy-stable distributions are utilized to characterize the data and determine the transverse-mass and centrality dependence of the Lévy parameters.
The Lévy parameterization is found to give a statistically acceptable description of the data with the Lévy exponent, $\alpha$, significantly greater than 1 and less than 2. This exponent is determined in 23 $m_{T}$ and in 6 centrality bins and is observed to be well described with its $m_{T}$-averaged value in every centrality bin. However, the transverse-mass-averaged values do depend on centrality.
The Lévy-scale parameter $R$ is proportional to the HWHM of the source distribution, with a coefficient of proportionality that depends on the Lévy exponent $\alpha$. The behavior of its inverse square as function of $m_{T}$ that is observed can be described by an affine linear fit, whose parameters are examined as a function of centrality. The parameter $A$, which can be related to the transverse velocity of the expansion, shows a trend similar to hydrodynamical predictions [39, 38]. However, the assumption of local thermalization results in the $\alpha=2$ special case (see, e.g., Refs. [39, 27, 38]), which is in significant contrast to our observations of $\alpha<2$ in each of the investigated centrality and transverse-mass bins.
Within statistical uncertainties, the parameter $B$, is consistent with zero or with a slightly negative value. A negative value of $B$ is possible if the Cooper-Frye freeze-out terms are also taken into account [62]. This result may indicate a source that includes local thermalization with hydrodynamical expansion, followed by rescattering and decays of resonances. However, such nonequilibrium, scale-dependent features typically would result in deviations from the Lévy shape and from the applicability of generalized central-limit theorems. But first-order deviations from the Lévy-stable source distributions using the expansion method of Refs. [25, 12, 13] were found to be consistent with zero. It is theoretically challenging to explain simultaneously the measured value of $\alpha$, which is found to be significantly less than the Gaussian value of 2, and the $m_{T}$ dependence of the Lévy scale parameter $R$, which follows a hydrodynamically predicted scaling (see, e.g., Ref. [40]).
@@
To quantify the suppression pattern, the Gaussian width and amplitude parameters, $\sigma$ and $H$, were introduced. These parameters are observed to be centrality independent, except in the 50%–60% case, where the uncertainty on $H$ increases and does not allow a statistically significant conclusion. This independence is helpful to rule out or validate models with predictions about this behavior. The pion laser model, for example, predicts a strong centrality and multiplicity dependence [63, 37, 43].
In Ref. [36] it was observed that the amount of suppression of $\lambda(m_{T})$ and hence of $\lambda(m_{T})/\lambda_{\rm max}$ at low $m_{T}$ can be an experimentally observable signal of the (partial) restoration of $U_{A}(1)$ symmetry and a measure of in-medium reduction of the mass of the anomalously heavy $\eta^{\prime}$ mass. An approximately constant trend of the intercept parameter $\lambda$ as a function of $m_{T}$ was observed in small systems at lower energies, ($E_{\rm LAB}=150$ AGeV [7, 20]), which is consistent with vanishing suppression ($H=0$). These results are both qualitatively and quantitatively different from the presence of the suppression ($H_{0}=0.42\pm 0.02$(stat) ), which is within uncertainties independent of the centrality in $\sqrt{s_{NN}}=200$ GeV Au$+$Au collisions at RHIC. This entirely experimental data-based, Monte-Carlo-independent observation suggests that the signal characterized by the value of $H$ is dependent on the energy and/or the system size. Further detailed measurements are needed to determine the energy and system size, where $H$ becomes nonvanishing for the first time, starting from zero in the small Be+Be collisions at the relatively low energy of $E_{\rm LAB}=150$ AGeV.
@@ Acknowledgements.
We thank the staff of the Collider-Accelerator and Physics Departments at Brookhaven National Laboratory and the staff of the other PHENIX participating institutions for their vital contributions. We acknowledge clarifying and inspiring discussions with Sándor Hegyi, Dubravko Klabučar, Robert Pisarski and György Wolf concerning the interpretation of their published results. S.L. thanks IFJ PAN, Krakow, Poland for their encouragement to finish this work. We acknowledge support from the Office of Nuclear Physics in the Office of Science of the Department of Energy, the National Science Foundation, Abilene Christian University Research Council, Research Foundation of SUNY, and Dean of the College of Arts and Sciences, Vanderbilt University (U.S.A), Ministry of Education, Culture, Sports, Science, and Technology and the Japan Society for the Promotion of Science (Japan), Conselho Nacional de Desenvolvimento Científico e Tecnológico and Fundação de Amparo à Pesquisa do Estado de São Paulo (Brazil), Natural Science Foundation of China (People’s Republic of China), Croatian Science Foundation and Ministry of Science and Education (Croatia), Ministry of Education, Youth and Sports (Czech Republic), Centre National de la Recherche Scientifique, Commissariat à l’Énergie Atomique, and Institut National de Physique Nucléaire et de Physique des Particules (France), J. Bolyai Research Scholarship, EFOP, HUN-REN ATOMKI, NKFIH, MATE KKP and OTKA (Hungary), Department of Atomic Energy and Department of Science and Technology (India), Israel Science Foundation (Israel), Basic Science Research and SRC(CENuM) Programs through NRF funded by the Ministry of Education and the Ministry of Science and ICT (Korea). Ministry of Education and Science, Russian Academy of Sciences, Federal Agency of Atomic Energy (Russia), VR and Wallenberg Foundation (Sweden), University of Zambia, the Government of the Republic of Zambia (Zambia), the U.S. Civilian Research and Development Foundation for the Independent States of the Former Soviet Union, the Hungarian American Enterprise Scholarship Fund, the US-Hungarian Fulbright Foundation, and the US-Israel Binational Science Foundation.
@@
Monte-Carlo simulations with scenarios that allow for (or exclude) a possible in-medium mass modification of the $\eta^{\prime}$ meson and their comparisons to data were performed in Refs. [36, 43, 44] and in the previous PHENIX paper (Ref. [1]) by utilizing standard $\chi^{2}$ and CL maps. Similar, but now centrality-dependent simulations are summarized in this Appendix. These comparisons result in $\chi^{2}$ and CL maps that are utilized to determine the $\chi^{2}$ minimum (or CL maximum) yielding the best values, as well as the statistical and the systematic uncertainties of a possible in-medium mass-drop of the $\eta^{\prime}$ meson. Figure 14 shows an example of such a comparison, where each bin corresponds to a comparison of Monte-Carlo simulations of resonance decay chains with our data on $\lambda(m_{T})/\lambda_{{\rm max}}$ as a function of centrality, as detailed in Section VII.
Each panel of Fig. 14 contains the optimal values of $m^{*}_{\eta^{\prime}}$ and $B_{\eta^{\prime}}^{-1}$ in the scanned region. For our simulations, a unique $\chi^{2}$ minimum is found, in contrast to earlier studies in Refs. [43, 44]. By now, it is well known, that the dominant mechanism for soft-particle production is a thermal one in $\sqrt{s_{{}_{NN}}}$= 200 GeV Au+Au collisions.
Earlier, a significantly broader class of models was considered in Refs. [43, 44] that resulted in two, characteristically different minima in the 0%–30% centrality class of the same reaction. Accordingly, Refs. [43, 44] evaluated the exclusion limits, focusing in particular on the smallest possible in-medium mass drop of the $\eta^{\prime}$ meson, which is required to describe the data with at least 0.1% CL in any of the considered Monte-Carlo simulations.
Our systematic investigations include an estimation of the exclusion limits for the lowest- and largest-possible values of the in-medium mass of the $\eta^{\prime}$ meson. The systematic variations are consistent, i.e. all the systematic variations result in essentially the same minimum. Thus, in contrast to Refs. [43, 44], our indirect observation of in-medium $\eta^{\prime}$ mass modification is used with the best value and its statistical and systematic uncertainties of $m^{*}_{\eta^{\prime}}$, rather than exclusion limits.
Table 2 shows the centrality-dependent, fitted values of the in-medium $\eta^{\prime}$ mass $m^{*}_{\eta^{\prime}}$ and the inverse slope $B^{-1}_{\eta^{\prime}}$ with statistical and systematic uncertainties. The exclusion limits of these parameters were evaluated, but they do not result in statistically significant new ranges.
@@
The $\eta^{\prime}$ spectrum has six parameters, which are considered as inputs to the $\lambda(m_{T})$ simulations. The two most important inputs are the in-medium mass of the $\eta^{\prime}$, $m^{*}_{\eta^{\prime}}$, and the characteristic-slope parameter, $B_{\eta^{\prime}}^{-1}$, of the $\eta^{\prime}$ mesons that are emitted with PDG mass after the decay of the condensate. The second parameter characterizes a second component, a low-$p_{T}$ part of the $\eta^{\prime}$ spectrum and determines the slope of the suppression of the $\lambda$ parameter at low-$p_{T}$ [36, 43, 44]. These two parameters determine the shape of the observed dip in the low-$m_{T}$ region of the $\lambda(m_{T})$ function: a strongly reduced $\eta^{\prime}$ mass causes a dip in the low-$m_{T}$ part of the $\lambda(m_{T})$ function, while $B_{\eta^{\prime}}^{-1}$ controls the slope of this dip, see Refs. [43, 44]. These two parameters are considered as fit variables. The values of four nuisance parameters $\alpha_{\rm th}$, $T_{f}$, $\langle u_{T}\rangle$, and $T_{\rm cond}$, are treated as constants in the fit, but are varied to evaluate the systematic uncertainties of the fitted values.
The $\chi^{2}$ scans provide fine grids of CL in each centrality class. The in-medium mass is scanned with 10 MeV steps between 0 and 958 MeV (the vacuum mass of $\eta^{\prime}$) and $B_{\eta^{\prime}}^{-1}$ with 20 MeV steps between 0 and 360 MeV. The minimum value of $\chi^{2}$ (or maximum value of CL) is determined, which provides the most likely values of $m^{*}_{\eta^{\prime}}$ and $B_{\eta^{\prime}}^{-1}$ and their statistical uncertainties, as well as the CL of the fits to the data, which are compared to the data in Fig. 11. The values of the four “constant” parameters, $\alpha_{{\rm th}}$, $T_{f}$, $\langle u_{T}\rangle$, and $T_{\rm cond}$, are varied to evaluate the systematic uncertainties of the fitted values:
$\alpha_{{\rm th}}$